OpenStudy (anonymous):

the equation formed by decreasing each root of ax^2 + bx + c = 0 by 1 is 2x^2 + 8x + 2 = 0. Then a. c=-a b. b=-c

OpenStudy (anonymous):

roots of this are \[-2\pm\sqrt{3}\] so roots of original equation are \[-1\pm\sqrt{3}\]

OpenStudy (anonymous):

1) Find roots of 2x^2 + 8x + 2 = 0 2) Define a function that has the same roots, only this time increased by one 3) Check which statement is true.

OpenStudy (anonymous):

i dint get both ...got something else!!

OpenStudy (anonymous):

multiply \[(x-(-1+\sqrt{3}))(x-(-1-\sqrt{3}))\] to get original equation

OpenStudy (anonymous):

which is actually a lot easier than it looks

OpenStudy (anonymous):

"first" \[xx=x^2\] "last" \[(-1+\sqrt{3})(-1-\sqrt{3})=1-3=-2\]

OpenStudy (anonymous):

x^2 -2x√3+2x+4 ??

OpenStudy (anonymous):

and "outer, inner" is \[(1+\sqrt{x})x+(1-\sqrt{3})x=2x\]

OpenStudy (anonymous):

to give \[x^2-2x+2=0\]

OpenStudy (anonymous):

i looks hard at first glance but it is easy to multiply this out. check what i wrote

OpenStudy (anonymous):

thanks a lot... :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

is the answer b ?